# First-Order Sparse Convex Optimization: Better Rates with Sparse Updates

### v2

@@ Appendix B Proof of Lemma 2
Denote $k=\textrm{rank}({\mathbf{X}})$. Let ${\mathbf{P}}\in\mathbb{S}^{n}$ denote the projection matrix onto the subspace $\textrm{Im}({\mathbf{X}})$ and let ${\mathbf{P}}_{\perp}\in\mathbb{S}^{n}$ denote the projection matrix onto the orthogonal complement of $\textrm{Im}({\mathbf{X}})$. Throughout the proof we let $\|{\cdot}\|_{1}$ and $\|{\cdot}\|_{2}$ denote the trace norm (sum of singular values) and the Euclidean (Frobenius) norm in some space of real (not necessarily symmetric) matrices, respectively. We begin by recalling two important facts. First, for any real (not necessarily symmetric) matrix ${\mathbf{A}}$ it holds that
$\|{{\mathbf{A}}}\|_{1}\leq\sqrt{\textrm{rank}({\mathbf{A}})}\|{{\mathbf{A}}}\|_{2}.$ (9)
Second, for any two real (not necessarily symmetric) matrices of appropriate dimension (such that the multiplication is defined) ${\mathbf{A}},{\mathbf{B}}$ it holds that
$\|{{\mathbf{A}}{\mathbf{B}}}\|_{2}\leq\|{{\mathbf{A}}}\|_{\infty}\|{{\mathbf{B}}}\|_{2},$ (10)
where we let $\|{{\mathbf{A}}}\|_{\infty}$ denote the largest singular value of ${\mathbf{A}}$.
Let us denote ${\mathbf{D}}={\mathbf{X}}-{\mathbf{Y}}$. It holds that,
$\|{{\mathbf{D}}}\|_{1}$ $=\|{({\mathbf{P}}+{\mathbf{P}}_{\perp}){\mathbf{D}}({\mathbf{P}}+{\mathbf{P}}_{\perp})}\|_{1}$
$\leq\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}+\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}+2\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}.$ (11)
We continue by upper-bounding each of the three terms in the RHS separately.
Note that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}})\leq k$. Thus, using Eq. (9) and then Eq. (10), and noting that $\|{{\mathbf{P}}}\|_{\infty}=1$, we have that
$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (12)
To bound the second term in the RHS of (B) we have,
$\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}$ $=\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}}\|_{1}\underset{(a)}{=}\textrm{Tr}({\mathbf{P}}_{\perp}{\mathbf{Y}})=\textrm{Tr}(({\mathbf{I}}-{\mathbf{P}}){\mathbf{Y}})$
$\underset{(b)}{\leq}\textrm{Tr}({\mathbf{X}})-\textrm{Tr}({\mathbf{P}}{\mathbf{Y}})=\textrm{Tr}({\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}})\leq\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}}}\|_{1}$
$\underset{(c)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2},$ (13)
where (a) holds since ${\mathbf{Y}}$ is positive semidefinite, and thus so is ${\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}$, (b) holds by the assumption that $\textrm{Tr}({\mathbf{Y}})=\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}=\textrm{Tr}({\mathbf{X}})$, and (c) is due to Eq. (12)
Finally, for the last term in the RHS of (B), noting that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp})\leq k$, using Eq. (9) and then Eq. (10) (and noting again that $\|{{\mathbf{P}}}\|_{\infty}=\|{{\mathbf{P}}_{\perp}}\|_{\infty}=1$), we have that
$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (14)
The lemma follows from plugging-in (12), (B), and (14) into the RHS of (B). ∎
Appendix C Sparse Projections onto the Bounded Positive Semidefinite Cone
@@
Let ${\mathbf{Z}}\in\mathbb{S}^{n}$ with eigen-decomposition ${\mathbf{Z}}=\sum_{i=1}^{n}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ ($\lambda_{1}\geq\lambda_{2}\geq...$) and $s\in[n]$. A Euclidean projection of ${\mathbf{X}}$ onto $\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R,\textrm{rank}({\mathbf{X}})\leq s\}$ is given by the matrix ${\mathbf{Z}}^{\prime}=\sum_{i=1}^{s}\lambda_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, where $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$ is the Euclidean projection of the vector $(\lambda_{1},\dots,\lambda_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
@@
[…]
@@
$=\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\lambda_{j}^{\prime 2}-2\langle{{\mathbf{Y}},{\mathbf{Z}}}\rangle+2\sum_{i=1}^{s}\lambda_{i}^{\prime}\lambda_{i}$
$\geq\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\lambda_{j}^{\prime 2}-2\sum_{i=1}^{s}\gamma_{i}\lambda_{i}+2\sum_{i=1}^{s}\lambda_{i}^{\prime}\lambda_{i},$
where the last inequality is due to von Neumann’s trace inequality.
@@
$\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}\geq\sum_{i=1}^{s}(\gamma_{i}-\lambda_{i})^{2}-\sum_{j=1}^{s}(\lambda_{j}^{\prime}-\lambda_{j})^{2}\geq 0,$
where the last inequality follows since $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$ is the Euclidean projection of $(\lambda_{1},\dots\lambda_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$. ∎

### v3

@@ Appendix D Proof of Lemma 2
Throughout the proof we let $\|{\cdot}\|_{1}$, $\|{\cdot}\|_{2}$, $\|{\cdot}\|_{\mathrm{op}}$ denote the nuclear norm, the Euclidean (Frobenius) norm, and operator norm (largest singular value) in $\mathbb{R}^{m\times n}$, respectively. Denote ${\mathbf{D}}:={\mathbf{Y}}-{\mathbf{X}}$ and $r=\mathrm{rank}({\mathbf{X}})$. Let ${\mathbf{X}}={\mathbf{U}}{\mathbf{S}}{\mathbf{V}}^{\top}$ be a compact SVD of $\mathbf{X}$}. Define the orthogonal projectors ${\mathbf{P}}_{{\mathbf{U}}}:={\mathbf{U}}{\mathbf{U}}^{\top},{\mathbf{P}}_{{\mathbf{V}}}:={\mathbf{V}}{\mathbf{V}}^{\top}$, and accordingly define the linear maps $\mathcal{P}_{T},\mathcal{P}_{\perp}:\mathbb{R}^{m\times n}\to\mathbb{R}^{m\times n}$ as
$\mathcal{P}_{T}({\mathbf{Z}}):={\mathbf{P}}_{{\mathbf{U}}}{\mathbf{Z}}+{\mathbf{Z}}{\mathbf{P}}_{{\mathbf{V}}}-{\mathbf{P}}_{{\mathbf{U}}}{\mathbf{Z}}{\mathbf{P}}_{{\mathbf{V}}},\qquad\mathcal{P}_{\perp}({\mathbf{Z}}):=({\mathbf{I}}-{\mathbf{P}}_{{\mathbf{U}}})\,{\mathbf{Z}}\,({\mathbf{I}}-{\mathbf{P}}_{{\mathbf{V}}}).$
Write
${\mathbf{D}}_{T}:=\mathcal{P}_{T}({\mathbf{D}}),\qquad{\mathbf{D}}_{\perp}:=\mathcal{P}_{T^{\perp}}({\mathbf{D}}),$
so that ${\mathbf{D}}={\mathbf{D}}_{T}+{\mathbf{D}}_{\perp}$.
Clearly, we have that
$\|{{\mathbf{D}}}\|_{1}\leq\|{{\mathbf{D}}_{T}}\|_{1}+\|{{\mathbf{D}}_{\perp}}\|_{1}.$ (10)
Additionally, since ${\mathbf{U}},{\mathbf{V}}$ are of rank $r$, it follows from the definition of the map $\mathcal{P}_{T}$ that $\mathrm{rank}({\mathbf{D}}_{T})\leq 2r$. Thus, we have that
$\|{{\mathbf{D}}_{T}}\|_{1}\leq\sqrt{\mathrm{rank}({\mathbf{D}}_{T})}\|{{\mathbf{D}}_{T}}\|_{2}\leq\sqrt{2r}\|{{\mathbf{D}}_{T}}\|_{2}.$ (11)
Thus, given (10) and (11), to prove the lemma, it suffices to show that I. $\|{{\mathbf{D}}_{\perp}}\|_{1}\leq\|{{\mathbf{D}}_{T}}\|_{1}$ and II. that $\|{{\mathbf{D}}_{T}}\|_{2}\leq\|{{\mathbf{D}}}\|_{2}$.
To see why $\|{{\mathbf{D}}_{T}}\|_{2}\leq\|{{\mathbf{D}}}\|_{2}$ holds, it suffices to show that $\mathcal{P}_{T}$ is an orthogonal projection, meaning it is a contraction in Frobenius norm. To see this, observe that by definition, for any $\mathbf{Z}$}, $\mathcal{P}_{\perp}(\mathcal{P}_{\perp}({\mathbf{Z}}))=\mathcal{P}_{\perp}({\mathbf{Z}})$, meaning $\mathcal{P}_{\perp}$ is an orthogonal projection. Since, by definition $\mathcal{P}_{T}=\mathcal{I}-\mathcal{P}_{\perp}$, where $\mathcal{I}$ is the identity map, it follows that $\mathcal{P}_{T}$ is also an orthogonal projection.
Thus, in the remainder of the proof we show that $\|{{\mathbf{D}}_{\perp}}\|_{1}\leq\|{{\mathbf{D}}_{T}}\|_{1}$.
Write the compact SVD of ${\mathbf{D}}_{\perp}=\widetilde{\mathbf{U}}\,\widetilde{\mathbf{S}}\,\widetilde{\mathbf{V}}^{\top}$. Since ${\mathbf{D}}_{\perp}=({\mathbf{I}}-{\mathbf{P}}_{{\mathbf{U}}}){\mathbf{D}}({\mathbf{I}}-{\mathbf{P}}_{{\mathbf{V}}})$, we have that
${\mathbf{U}}^{\top}\widetilde{\mathbf{U}}=\mathbf{0},\qquad{\mathbf{V}}^{\top}\widetilde{\mathbf{V}}=\mathbf{0}.$ (12)
Define ${\mathbf{W}}:=\widetilde{\mathbf{U}}\widetilde{\mathbf{V}}^{\top}$. Note that $\|{\mathbf{W}}\|_{\mathrm{op}}=1$. Note that due to (12) we have that $\|{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}}}\|_{\mathrm{op}}=1$.
Using the variation characterization of the nuclear norm, we have that
$\|{{\mathbf{Y}}}\|_{1}$ $=\|{{\mathbf{X}}+{\mathbf{D}}}\|_{1}$
$=\max_{\|{{\mathbf{Q}}}\|_{\mathrm{op}}\leq 1}\langle{{\mathbf{Q}},{\mathbf{X}}+{\mathbf{D}}}\rangle$
$\underset{(a)}{\geq}\langle{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}},{\mathbf{X}}+{\mathbf{D}}}\rangle$
$=\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{X}}}\rangle+\langle{{\mathbf{W}},{\mathbf{X}}}\rangle+\langle{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}},{\mathbf{D}}}\rangle$
$\underset{(b)}{=}\|{{\mathbf{X}}}\|_{1}+\langle{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}},{\mathbf{D}}_{T}+{\mathbf{D}}_{\perp}}\rangle$
$\underset{(c)}{=}\|{{\mathbf{X}}}\|_{1}+\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{D}}_{T}}\rangle+\langle{{\mathbf{W}},{\mathbf{D}}_{\perp}}\rangle$
$\underset{(d)}{=}\|{{\mathbf{X}}}\|_{1}+\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{D}}_{T}}\rangle+\|{{\mathbf{D}}_{\perp}}\|_{1},$
where (a) holds since, as noted above, $\|{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}}}\|_{\mathrm{op}}=1$, (b) holds since using the SVD of $\mathbf{X}$} we have that $\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{X}}}\rangle=\mathrm{Tr}({\mathbf{U}}{\mathbf{V}}^{\top}{\mathbf{X}}^{\top})=\mathrm{Tr}({\mathbf{U}}{\mathbf{V}}^{\top}{\mathbf{V}}{\mathbf{S}}{\mathbf{U}}^{\top})=\mathrm{Tr}({\mathbf{S}})=\|{{\mathbf{X}}}\|_{1}$, (c) holds since by definition of $\mathbf{D}$$_{\perp}$ we have that $\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{D}}_{\perp}}\rangle=0$ and by definition of ${\mathbf{D}}_{T}$ and $\mathbf{W}$} we have that $\langle{{\mathbf{W}},{\mathbf{D}}_{T}}\rangle=0$, and (d) holds since, by definition of $\mathbf{W}$} and using the SVD of $\mathbf{D}$$_{\perp}$, we have that $\langle{{\mathbf{W}},{\mathbf{D}}_{\perp}}\rangle=\mathrm{Tr}(\widetilde{\mathbf{U}}\widetilde{\mathbf{V}}^{\top}{\mathbf{D}}_{\perp}^{\top})=\mathrm{Tr}(\widetilde{\mathbf{U}}\widetilde{\mathbf{V}}^{\top}\widetilde{\mathbf{V}}\widetilde{\mathbf{S}}\widetilde{\mathbf{U}}^{\top})=\mathrm{Tr}(\widetilde{\mathbf{S}})=\|{{\mathbf{D}}_{\perp}}\|_{1}$.
Using the assumption that $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$ and the fact that $\|{{\mathbf{U}}{\mathbf{V}}^{\top}}\|_{\mathrm{op}}=1$, we indeed have that
$\|{{\mathbf{D}}_{\perp}}\|_{1}\leq-\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{D}}_{T}}\rangle\leq\|{{\mathbf{U}}{\mathbf{V}}^{\top}}\|_{\mathrm{op}}\cdot\|{{\mathbf{D}}_{T}}\|_{1}=\|{{\mathbf{D}}_{T}}\|_{1},$
which concludes the proof. ∎
Appendix E Sparse Projections onto the Nuclear Norm Ball
@@
Let ${\mathbf{Z}}\in\mathbb{R}^{m\times n}$ with SVD ${\mathbf{Z}}=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}{\mathbf{u}}_{i}{\mathbf{v}}_{i}^{\top}$ ($\sigma_{1}\geq\sigma_{2}\geq$...) and $s\in[\min\{m,n\}]$. A Euclidean projection of $\mathbf{Z}$} onto $\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\|{{\mathbf{X}}}\|_{1}\leq R,\mathrm{rank}({\mathbf{X}})\leq s\}$ is given by ${\mathbf{Z}}^{\prime}=\sum_{i=1}^{s}\sigma_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{v}}_{i}^{\top}$, where $(\sigma_{1}^{\prime},\dots,\sigma_{s}^{\prime})$ is the Euclidean projection of the vector $(\sigma_{1},\dots,\sigma_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
@@
Given some ${\mathbf{Z}}\in\mathbb{R}^{m\times n}$ and $s\in[\min\{m,n\}]$, let $\mathbf{Z}$$^{\prime}$ be as described in the lemma, and let ${\mathbf{Y}}\in\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\|{{\mathbf{X}}}\|_{1}\leq R,~\mathrm{rank}({\mathbf{X}})\leq s\}$ given by its SVD ${\mathbf{Y}}=\sum_{i=1}^{s}\gamma_{i}{\mathbf{p}}_{i}{\mathbf{w}}_{i}^{\top}$, which clearly implies that $(\gamma_{1},\dots,\gamma_{s})\in\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$. Recall $\|{\cdot}\|$ denotes the Euclidean norm. It holds that
@@
$=\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\sigma_{j}^{\prime 2}-2\langle{{\mathbf{Y}},{\mathbf{Z}}}\rangle+2\sum_{i=1}^{s}\sigma_{i}^{\prime}\sigma_{i}$
$\geq\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\sigma_{j}^{\prime 2}-2\sum_{i=1}^{s}\gamma_{i}\sigma_{i}+2\sum_{i=1}^{s}\sigma_{i}^{\prime}\sigma_{i},$
where the last inequality is due to von Neumann’s trace inequality [21].
@@
$\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}\geq\sum_{i=1}^{s}(\gamma_{i}-\sigma_{i})^{2}-\sum_{j=1}^{s}(\sigma^{\prime}_{j}-\sigma_{j})^{2}\geq 0,$
where the last inequality holds since $(\sigma_{1}^{\prime},\dots,\sigma_{s}^{\prime})$ is the Euclidean projection of $(\sigma_{1},\dots\sigma_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$. ∎